### v1

# Localizing Factual Inconsistencies in Attributable Text Generation

Abstract
There has been an increasing interest in detecting hallucinations in model-generated texts, both manually and automatically, at varying levels of granularity. However, most existing methods fail to precisely pinpoint the errors. In this work, we introduce QASemConsistency, a new formalism for localizing factual inconsistencies in attributable text generation, at a fine-grained level. Drawing inspiration from Neo-Davidsonian formal semantics, we propose decomposing the generated text into minimal predicate-argument level propositions, expressed as simple question-answer (QA) pairs, and assess whether each individual QA pair is supported by a trusted reference text. As each QA pair corresponds to a single semantic relation between a predicate and an argument, QASemConsistency effectively localizes the unsupported information. We first demonstrate the effectiveness of the QASemConsistency methodology for human annotation, by collecting crowdsourced annotations of granular consistency errors, while achieving a substantial inter-annotator agreement ($\kappa>0.7)$. Then, we implement several methods for automatically detecting localized factual inconsistencies, with both supervised entailment models and open-source LLMs. [^1] Our codebase, dataset, and models can be found at https://github.com/ariecattan/qasem_consistency
1 Introduction
Figure 1: An annotation example of a localized factual consistency error according to our QASemConsistency methodology. Here, the model incorrectly identifies the location of the paraglider’s fall as the location where the rescue team was training, indicated by the QA “Where did someone fall? in the Annalong Valley in Countly Antrim”.
Large Language Models (LLMs) are used very effectively across a broad range of text generation tasks. However, despite remarkable progress in recent years, LLMs remain prone to generating factual inconsistencies. This phenomenon, commonly referred to as “hallucinations”, limits their broader deployment and utility (Huang et al., 2023).
This work focuses on attributable text generation, where the generated content can be verified against a trusted supporting source, referred here as the “reference text”. This reference text may be part of the input for generation, as typical in text summarization or open book QA (Gao et al., 2023b; Slobodkin et al., 2024), retrieved post-generation (Bohnet et al., 2022; Gao et al., 2023a; Min et al., 2023), or identified by the models themselves when instructed to provide citations to external sources (Liu et al., 2023a; Yue et al., 2023).
To address hallucinations, there has been increasing research interest in identifying unsupported content in model-generated text, both manually and automatically. This task is typically framed as a textual entailment problem, requiring that the generated text should be supported (entailed) by the reference text. Detecting unsupported information is valuable for multiple purposes. For evaluating the factual consistency of attributable text generation, both human annotation protocols and automated inconsistency detection models are needed (Honovich et al., 2022; Gekhman et al., 2023). Automated inconsistency detection can further provide valuable feedback for end users about suspected unsupported content in the LLMs’ output, and also contributes to model improvements, via post-editing (Gao et al., 2023a), enhanced training (Nan et al., 2021; Wan and Bansal, 2022; Roit et al., 2023), self-critique (Wadhwa et al., 2024), or by imposing constraints during decoding (Wan et al., 2023).
To better fulfil these goals, it is desired to pinpoint which parts of the generated text are not supported, especially as LLMs continue to improve such that factual inconsistencies become more localized, as illustrated in Figure 1. While recent research has made useful strides in finer-grained inconsistency detection, ranging from entire texts to sentences, claims, and even question-generation and question-answering based solutions, these sub-sentence representations often remain insufficiently granular. For example, an “atomic” claim in FActScore Min et al. (2023) is still based on multiple predicate-argument relations, and each of them might be either supported or unsupported (see Table 1, §2.1 and §6).
In this work, we introduce QASemConsistency, a novel protocol for detecting localized factual inconsistencies in attributable text generation, applicable to both human and automatic detection. Inspired by Neo-Davidsonian formal semantics, our method decomposes the generated text into elementary assertions, in the form of atomic question-answer (QA) pairs (QASRL (He et al., 2015) and QANom (Klein et al., 2020)), where each pair corresponds to a single predicate-argument relation. Localizing factual inconsistencies then involves identifying the set of QA pairs that are not supported by the reference text. Figure 1 illustrates our QASemConsistency methodology by representing the factual inconsistency via a simple QA “Where did someone fall? in the Annalong Valley in County Antrim”, which is not supported by the reference text. By assessing each individual predicate-argument level statement, QASemConsistency can pinpoint more precisely to the factual mistakes than prior methods (illustrated in Table 1).
Furthermore, since we represent predicate-argument relations with simple natural language expressions (questions and answers), our QASemConsistency methodology is well-suited for manually annotating granular consistency errors. Indeed, we collect the first dataset with localized annotations of factual inconsistency at the predicate-argument level via crowdsourcing and achieve a high inter-annotator agreement (> 0.7 Fleiss Kappa) (§4).
Finally, we implement methods for automatically detecting whether each individual QASem QA is supported by the reference text, following the QASemConsistency methodology (§5). We conduct experiments with a variety of models, including supervised NLI models and prompting several LLMs. While these models were supervised on standard entailment datasets, our results show that they can effectively handle our fine-grained QA annotation, providing more detailed error detection.
2 Background
2.1 Fine-grained Detection of Factual Inconsistencies
Source Gareth Colfer-Williams, 25, died last week at his home in Swansea, the city at the centre of an epidemic of the disease which has reached 942 cases. But the examination was unable to establish whether measles was the main cause of his death. An inquest will be opened and adjourned on Tuesday to allow further tests… Public Health Wales said on Friday that laboratory tests confirmed a diagnosis of measles but further tests were needed to determine the cause of death…
Summary An inquest into the death of a man who died of measles has been opened and adjourned after a post-mortem examination failed to establish how he got the illness.
Token-level (CLIFF) An inquest into the death of a man who died of measles has been opened and adjourned after a post-mortem examination failed to establish how he got the illness.
QGQA ($Q^{2}$) - An inquest: What has been opened and adjourned into the death of a man who died of measles? - A man: Who died of measles? - measles: What was the cause of death of a man? - a post-mortem examination: What failed to establish how he contracted measles? - the illness: A post-mortem examination failed to establish how he got what?
Claim-level (FActScore) - An inquest into the death of a man has been opened. - The man died of measles. - A post-mortem examination was conducted. - The post-mortem examination failed to establish how he got measles. - The inquest has been adjourned.
QASemConsistency died/death: - Who died? A man - How someone died? from measles failed: - What failed to do something? a post-mortem examination - What did something fail to do? to establish how he got the illness got: - Who got something? he; A man - What did someone get? the illness; measles opened: - What has been opened? An inquest into the death of a man establish: - What didn’t establish something? a post-mortem examination - What didn’t something establish? how he got the illness examination: - Who was examined? A man who died of measles - When was someone examined? post-mortem - Why was someone examined? to establish how he got the illness
Table 1: Comparison of QASemConsistency to existing fine-grained decompositions. Here, the summary is inconsistent in multiple respects: it assumes that the man died from measles, while the cause of death remains unclear; the examination actually failed to establish the cause of death rather than how the man got the illness; and an inquest will be opened in the future rather than in the past. The token-level annotations are from CLIFF (Cao and Wang, 2021), with highlighted tokens indicating those marked as expressing unsupported information. For QGQA and claim-level, indicates “supported” and indicates “not supported”. QGQA lists the candidate answers (noun phrases) and the generated questions using $Q^{2}$ (Honovich et al., 2021), where the consistency labels indicate whether the candidate answer from the summary corresponds to the predicted answer based on the source. FActScore claims are generated by GPT 4o given FActScore’s instructions and demonstrations (Min et al., 2023), with our annotation of factual consistency. Unlike the other methods, our QASemConsistency methodology decomposes the summary into fine-grained predicate-argument statements in the form of QAs, pinpointing more precisely at the unsupported facts. For example the unsupported QA “How someone died? from measles” pinpoints the inconsistency about the cause of death, while FActScore mixes in the same claim both the cause of death as well as the indication of who died, where the latter is consistent in the summary.
Current efforts in detecting factual consistency errors continuously progress towards more localized methods that pinpoint the inconsistent information. Starting from approaches that highlight individual sentences (Laban et al., 2022), inspect “facts” (Min et al., 2023), or use question-generation and question-answering (Honovich et al., 2021; Fabbri et al., 2022) schemes to decompose the information in the generated text. While these decomposition techniques result in intuitive propositions in natural language, they often lack granularity and include multiple units of information, each of which might be supported or not by the reference text. As exemplified in Table 1 (Claim-level), the extracted claim “the man died of measles” could be further decomposed into two semantic relations - (1) “the man died” (PATIENT) and (2) “the death was due to measles” (CAUSE), with only the latter being unsupported by the reference text. We provide a detailed review of the relevant literature in Section 6.
Conversely, detecting factual inconsistencies at a finer level of granularity has involved using complex syntactic (Goyal and Durrett, 2020) and semantic (Ribeiro et al., 2022) formalisms. These formal approaches require linguistic expertise from annotators to manually evaluate generative models. Consequently, these methods were applied only for automatic detection of factual inconsistencies with models trained on synthetic data.
In this work, we represent minimal propositions, which correspond to individual predicate-argument relations, using intuitively comprehensible natural language question-answer pairs. This enables both human and automatic detection and a fine-grained representation of factual consistency errors.
2.2 Predicate-argument Relations: From Neo-Davidsonian Semantics to QA-SRL
Neo-Davidsonian semantics (Davidson, 1967; Higginbotham, 1983; Parsons, 1990) is a framework for representing the complex interactions between events and participants (i.e. predicate-argument relations) in a logical form. For example, given the sentence “Mary bought a book yesterday and gave it to John with a smile”, the Neo-Davidsonian representation is:
{dmath}
∃e_1, e_2   (Buying(e_1) ∧Agent(e_1, Mary) ∧Theme(e_1, Book) ∧Time(e_1, Yesterday) ∧Giving(e_2) ∧Agent(e_2, Mary) ∧Recipient(e_2, John) ∧Theme(e_2, Book) ∧Manner(e_2, Smile))
This representation indicates that there are two predicates — buy $(e_{1})$ and give $(e_{2})$, each with its own set of arguments.
Following the underlying principles of Neo-Davidsonian semantics, various approaches have been developed to model fine-grained propositions corresponding to individual predicate-argument relations, including FrameNet (Baker et al., 1998), PropBank (Palmer et al., 2005), Semantic Dependency Parsing (Oepen et al., 2014) and AMR (Banarescu et al., 2013). These approaches, however, typically rely on complex semantic formalisms, making them less accessible to non-expert annotators, and in a sense harder to extract and manipulate with LLMs.
To bridge this gap, we propose using QA-SRL, a semantic formalism that simplifies over traditional SRL schemes by representing each predicate-argument relation through a simple and “minimal” question-answer pair, such as “Who bought something? Mary” (He et al., 2015; FitzGerald et al., 2018; Roit et al., 2020). For example, the QA-SRL representation of the sentence above is detailed in 2. Here, “Mary” is identified as the agent of the predicate “bought”, where this semantic relation is represented by the question “Who bought something? Mary”. Loosely speaking, each QA typically corresponds to a single Neo-Davidsonian proposition.
bought Who bought something? Mary
What did someone buy? A book
When did someone buy something?
   Yesterday
gave Who gave something? Mary
Who gave something to? John
What did someone give? A book
How did someone give? With a smile
Table 2: QA-SRL representation for “Mary bought a book yesterday and gave it to John with a smile.”
By relying on a comprehensible natural-language representation, QA-SRL largely subsumes traditional SRL schemes. Notably, it successfully covers valuable implicit semantic arguments (Roit et al., 2024), which are intuitively captured by human annotators, and subsequently by models trained on such annotated data. The QA-SRL formalism has been expanded to include discourse relations (e.g., QADiscourse; Pyatkin et al., 2020), deverbal nominalizations (e.g., QANom; Klein et al., 2020) and adjectival predicates  (QA-Adj; Pesahov et al., 2023), along with the development of a unified QASem framework, covering multiple predicate types (Klein et al., 2022).
3 QASemConsistency
Given a generated text $\mathcal{Y}$ and a reference text $\mathcal{X}$ that is expected to support the information in $\mathcal{Y}$ (e.g. the source for generation or a grounding text for it), we define the task of localizing factual inconsistencies as identifying the set of “localized” assertions that are contained in $\mathcal{Y}$ but are not supported by $\mathcal{X}$. This involves the decomposition of $\mathcal{Y}$ into such assertions, each being a unit of information that can be individually assessed for its entailment by $\mathcal{X}$. For effective localization, we suggest two desired properties of this decomposition. First, the assertions should be as minimal in scope as possible, where ideally each assertion should not be further decomposable into smaller verifiable assertions. Second, each assertion should be human interpretable, providing clear insights to common language speakers, thus allowing efficient crowdsourced annotation for localized factual inconsistencies.
To fulfill these properties, we propose decomposing $\mathcal{Y}$ into the set of its predicate-argument level propositions, using the QASem framework (§2.2). By construction, each question-answer pair (QA) in QASem corresponds to a single predicate-argument relation, expressed in natural language.
[…]
Once all QAs are assigned with a label, we can also calculate an overall factual consistency score, defined as the percentage of supported QAs over all QAs. This interpretable score represents the proportion of supported semantic relations within $\mathcal{Y}$ and can be used for model evaluation. For example, the overall factual consistency score of the generated text in Table 1 is 7/12.222In practice, there are a few more QAs that we omitted for brevity. For comparison, the prior localization methods of QGQA and FActScore assign a score of 1/5 to the same summary. This discrepancy stems from the insufficient granularity level of these approaches (see §2.1 and §6), where each assertion contains a mixture of supported and unsupported information.
QASem parsing.
In this paper, we generate QAs for both verbal and nominal predicates using a parser that we purposely retrained on the QASRL and QANom datasets, improving over the original QASem parser (Klein et al., 2022). Such QAs are generated for each generated text $\mathcal{Y}$, and are then used to detect localized factual inconsistencies by both human annotation (§4) and automated models (§5). See training details and performance of our parser (to be released upon publication) in Appendix A.
4 Human Detection of Localized Factual Inconsistencies
In this section, we present the QASemConsistency evaluation methodology for manually annotating localized factual inconsistencies in generated texts supported by a corresponding annotation tool (Figure 2), and assess its effectiveness empirically with crowd annotators (§4.1). Our experiments demonstrate the high efficacy of our approach by achieving a high inter-annotator agreement, resulting in an annotation of granular consistency errors (§4.2). We further release a dataset of our annotations to foster future research and evaluation of factual-inconsistency detection models  (§4.3). To the best of our knowledge, this is the first work to annotate factual inconsistencies for predicate-argument level propositions.
4.1 Human Evaluation Process
As mentioned in (§3), given a generated text $\mathcal{Y}$, we predict QA pairs that represent localized propositions corresponding to single predicate-argument relations, using our novel QASem parser. [^3] In some cases, low-quality generated QAs were filtered in the annotation process, in a preliminary step (see Appendix B). Then, human annotators inspect these QAs along with the reference text $\mathcal{X}$ and the generated text $\mathcal{Y}$ and determine for each QA whether it is supported by $\mathcal{X}$ or not. Table 1 (QASemConsistency) shows an example of such annotations. We now describe our complete annotation process.
Figure 2: An example of the QA annotation step. The article in the left side is the reference text and the summary is shown on the right side. At each time, the interface highlights a specific predicate (here “died”) and displays the QAs representing the predicate-argument relations for that predicate. The green thumbs-up and red thumbs-down correspond to supported and not supported, respectively.
Enhancing annotation efficiency.
We first observe that if an entity mentioned in $\mathcal{Y}$ is absent in the reference text $\mathcal{X}$ (i.e., an extrinsic hallucination), all QA pairs featuring this entity as the answer are inevitably not supported. For instance, considering the source article in Table 1, if the summary would have mentioned “An inquest into the death of a woman…”, then the QAs “Who died? A woman” and “Who got something? A woman” would not be supported since the reference text does not mention a woman.
Leveraging this observation, we make the annotation process more efficient by dividing it into two sequential steps. In the first step, annotators go through the entity spans from the generated text $\mathcal{Y}$ corresponding to QASem arguments (i.e. the answers) and classify each span to either “covered” or “not covered”, according to whether it is mentioned explicitly or can be implied from $\mathcal{X}$. Any answer classified as “not covered” (representing an extrinsic hallucination) automatically renders all associated QA pairs as “unsupported”. This eliminates the need for annotators to evaluate these QA pairs individually.
In the second step, annotators focus exclusively on the remaining QA pairs – those whose answers have been confirmed to be covered by the reference text $\mathcal{X}$. We define a QA as “supported” if the meaning of that QA can be deduced from $\mathcal{X}$. To enhance annotators’ focus, all QAs of the same predicate are shown together. In addition to the “support” labels, annotators were encouraged to write free-text notes to justify their decisions, encouraging deeper consideration. These two separate stages improve annotation efficiency, and also introduce an additional layer that classifies factual inconsistencies into extrinsic versus intrinsic errors.
Annotation Tool.
To facilitate the human annotation process, we develop an intuitive annotation interface which streamlines the two steps. [^4] Our tool will be made publicly available. Figure 2 shows the interface of the second annotation step (QA evaluation). See Appendix B for implementation details.
Datasets and Generative Models.
We collect human annotations from two domains and generation settings. First, we consider the task of abstractive summarization on the XSUM dataset, that summarizes news article to a single sentence. Specifically, we sample a subset of 74 summaries from the CLIFF dataset (Cao and Wang, 2021), in which the source articles are from XSUM (Narayan et al., 2018) and the summaries were automatically generated by BART (Lewis et al., 2020) and PEGASUS (Zhang et al., 2020). CLIFF manually annotated each generated summary with token level annotation of consistency errors. Second, we annotate faithfulness localization for 34 generated biographies (3-5 sentences) included in FActScore (Min et al., 2023), which are generated by LLMs such as ChatGPT, InstructGPT, and the retrieval-augmented PerplexityAI model.555perplexity.ai FActScore manually labelled each extracted claim as supported or not against the Wikipedia page of the subject of the biography.
For both datasets, we collect annotations from 3 different workers for each generated text.
Annotators.
We recruit annotators through Amazon Mechanical Turk. [^6] https://www.mturk.com/ Each HIT consists of one reference text, $\mathcal{X}$, and two model outputs, $\mathcal{Y}_{1}$ and $\mathcal{Y}_{2}$. For XSUM, each HIT includes a source article with summaries generated by BART and PEGASUS, and compensated with \$1.50 per HIT. For FActScore, we select two model outputs for each instance, with a compensation of \$3.00 per HIT, reflecting the larger number of QAs to annotate.
Annotation Guidelines.
We follow the controlled crowdsourcing protocol (Roit et al., 2020), which consists of training the workers with detailed instructions (using intuitive slides) and providing ongoing personalized feedback throughout the process. Multiple examples were included to illustrate these two definitions. To further assist annotators with QA evaluation, we advise them to rephrase the question-answer pair as an affirmative statement and assess whether the reference text supports this statement. For instance, the QA pair “What did someone open? An investigation” could be rephrased as “Someone opened an investigation”.
Importantly, in line with our definition of factual consistency as an entailment task (§3), annotators were instructed to rely solely on the reference text. The use of external resources (e.g., web search) was restricted for clarifying the definitions of complex terms (e.g., “six-under-par”), but not to verify content not explicitly stated in the reference text.
4.2 Advantages of QASemConsistency
High Inter-Annotator Agreement.
We measure inter annotator agreement using Fleiss Kappa on all annotated QAs. We achieve a $\kappa$ = 0.72 for XSUM and $\kappa=0.79$ for FActScore, indicating a substantial agreement. For comparison, previous factuality benchmarks report lower agreement: Pagnoni et al. (2021) report $\kappa=0.58$ at the sentence-level and  Cao and Wang (2021) reports $\kappa=0.35$ at the token-level. Annotating factual consistency is a challenging and sometimes subjective task (Falke et al., 2019). We believe that the high agreement is due to the QASemConsistency’s decomposition into atomic semantic relations, where annotators need to assess a single assertion at a time. In addition, by collecting factual consistency annotation for each predicate-argument relation, QASemConsistency ensures that the human raters will assess all details.
QASemConsistency vs. Token-Level Annotations.
We compare the localized annotations of supported QAs in the CLIFF’s subset of our QASemConsistency dataset to the original token-level annotations in CLIFF. Specifically, we analyze the cases where summaries were annotated as fully supported in CLIFF but not in our annotations and vice-versa. Notably, we observe that all ten summaries where all QAs were annotated as supported in our dataset were also annotated as fully supported in CLIFF. Conversely, we identified 10 out of 74 summaries marked as fully supported in CLIFF whereas our annotators found unsupported QAs. Upon analyzing these cases, we discover that the majority of the summaries (7 out of 10) were indeed not fully supported by the reference text, as reflected by the unsupported QAs. For instance, the summary sentence “The US has suspended its participation in talks with Russia to try to broker a cessation of hostilities in Syria, the State Department says.” is not fully supported by the reference text “The United States is suspending its participation in bilateral channels with Russia that were established to sustain the cessation of hostilities” because the reference text mentions that the suspension of the U.S participation is “to sustain the cessation” and not “to broker a cessation”. This fine-grained factual inconsistency was indicated by several QAs marked as not supported, such as “Why has someone suspended something? To try to broker a cessation of hostilities in Syria” about the predicate “suspended”. Two other summaries could be interpreted as either supported or not supported, with our annotators showing localized disagreement on these cases. One summary was mistakenly annotated as not supported by our annotators. This analysis confirms that having human annotators verify each predicate-argument relation is not only beneficial for localizing factual inconsistencies, but also effectively helps annotators to identify them.
4.3 Evaluation Benchmark
Dataset #Summaries #QAs IAA
XSUM 74 692 0.72
FActScore 36 935 0.79
Table 3: Statistics of our collected benchmark QASemConsistency.
Overall, we gathered a dataset of entailment annotations for 1627 QASem QAs (see Table 3 for more statistics). The ground truth label for each QA is determined by the majority vote among the annotators. This dataset allows evaluating the performance of automatic detectors of localized factual inconsistencies, as we do in §5, and is released (upon publication) to facilitate future research of inconsistency detection models.
5 Automatic Detection of Localized Factual Inconsistencies
In this section we describe several models that automatically localize factual inconsistencies according to our methodology. We decompose the generated text $\mathcal{Y}$ into a list of QA pairs $\{qa_{1},qa_{2},...,qa_{n}\}$, and define the likelihood for each $qa_{i}$ to be supported (entailed) by the reference text $\mathcal{X}$ as $s(\mathcal{X},qa_{i})\in[0,1]$. We use the QASem parser from  Section 3 to generate the QAs and conduct our experiments with different entailment classifiers.
5.1 Entailment Classifiers
We experiment with two types of models:
Supervised
We apply two recent off-the-shelf NLI models to predict whether the QASem QA $qa_{i}$ is entailed by the reference text. In these experiments, the premise is the reference text $\mathcal{X}$ and the hypothesis is the concatenation of the question $q_{i}$ and the answer $a_{i}$ in $qa_{i}$.
The first model is DeBERTa-NLI, an encoder-only model based on DeBERTA-V3-Large (He et al., 2023), which was trained over a mixture of challenging NLI datasets, including MNLI (Williams et al., 2018), ANLI (Nie et al., 2020), FEVER (Thorne et al., 2018), WNLI (Liu et al., 2022) and LingNLI (Parrish et al., 2021). [^7] https://huggingface.co/MoritzLaurer/DeBERTa-v3-large-mnli-fever-anli-ling-wanli This model includes a classification head that classifies the input to entailment, neutral and contradiction. We consider the probability of the entailment class as our score $s(\mathcal{X},qa_{i})$.
The second supervised classifier is from TRUE (Honovich et al., 2022), an encoder-decoder model based on T5-XXL (Raffel et al., 2020) and finetuned on diverse entailment datasets: SNLI (Bowman et al., 2015), MNLI (Williams et al., 2018), FEVER (Thorne et al., 2018), Scitail (Khot et al., 2018), PAWS (Zhang et al., 2019) and VitaminC (Schuster et al., 2021). [^8] https://huggingface.co/google/t5_xxl_true_nli_mixture The TRUE model, also known as AutoAIS (Bohnet et al., 2022), was widely used in previous work for automatically measuring factual consistency (Bohnet et al., 2022; Gao et al., 2023a; Roit et al., 2023; Gao et al., 2023b; Slobodkin et al., 2024). This model was trained to generate “1” if the hypothesis is entailed by the premise and “0” otherwise. Hence, $s(\mathcal{X},qa_{i})$ is set as the probability of generating “1”, given the likelihood score of “1” and “0”.
LLM Prompting
We also use several LLMs to predict $s(\mathcal{X},qa_{i})$ using zero-shot prompting. [^9] We also experimented with few-shot demonstrations on a sample of the QASemConsistency dataset but these did not improve performance. We instruct the model to write “Yes” if the proposition corresponding to the QA is entailed by the reference text, and “No” otherwise. The prompt used in our experiments is presented in Appendix C. We apply this prompt to recent open-source LLMs including Llama 8B (v3.1) (Dubey et al., 2024), Gemma 7B (v1.1) (Team et al., 2024), Gemma 9B (v2) (Riviere et al., 2024), Mistral 7B (v0.2) (Jiang et al., 2023) and Mixtral 8x7B (v0.1) (Jiang et al., 2024). We do not use commercial models because our evaluation is based on the probability score $s(\mathcal{X},qa_{i})\in[0,1]$ of generating “Yes” or “No” (see §5.2). For all prompted models, we define the entailment score $s$ as the probability of generating “Yes” as the first generated token, based on the likelihood score of generating “Yes” and “No”.
5.2 Evaluation of Automatic Detection
We apply the above models to predict an entailment score $s(\mathcal{X},qa_{i})$ for all QAs in our collected QASemConsistency dataset and compare the prediction to the ground truth. We report two metrics to evaluate the performance of our automatic models in detecting localized factual inconsistencies. First, we follow standard evaluation practices (Honovich et al., 2022) and report the Area Under the Receiver Operating Characteristic Curve (ROC AUC), which plots the true positive rate against the false positive rate for different possible thresholds of $s(\mathcal{X},qa_{i})$. Second, since the models are trained or instructed to classify a QA as either supported or not supported, we also measure the performance of the hard prediction, considering a QA as supported if $s(\mathcal{X},qa_{i})\geq 0.5$. As our main goal is to localize unsupported QAs, which becomes more and more subtle in recent models, we propose to report the F1 score for identifying unsupported QAs. That is, we consider a true positive instance when both the ground truth and the model predict “not supported”.
5.3 Results
CLIFF FActScore
F1 AUC F1 AUC
Supervised NLI models
DeBERTa-NLI (304M) 37.2 67.9 47.0 75.9
TRUE (11B) 62.2 69.9 65.5 86.4
LLM prompting
Gemma v1.1 (7B) 52.4 72.7 71.3 84.1
Gemma v2 (9B) 53.7 78.3 76.6 90.2
Mistral v0.3 (7B) 43.9 71.7 67.8 84.1
Mixtral v0.1 (8x7B) 57.5 69.2 70.8 80.6
Llama v3.1 (7B) 47.6 74.3 72.4 88.9
Table 4: Results of automatic evaluators on our collected benchmark QASemConsistency.
Table 4 presents the results of our automatic evaluation. Notably, Gemma v2 (9B) achieves the highest ROC AUC scores, reaching 78.3 on CLIFF and 90.2 on FActScore. The TRUE model achieves the highest F1 score on CLIFF (62.2) and Gemma v2 (9B) achieves the highest F1 score on FActScore (76.6). These results suggest that even though these models were trained on standard entailment datasets, where the hypothesis is a single sentence, they can effectively generalize to the hypotheses in our setting, that is, QASem QAs that represent predicate-argument relations (e.g, “How someone died? from measles”). This ability is likely as such models were trained on a large amount of text while QASem QAs are also natural language expressions.
CLIFF FActScore
Prec. Rec. Prec. Rec.
Supervised NLI models
DeBERTa-NLI (304M) 73.1 25.8 73.7 36.4
TRUE (11B) 46.7 93.2 51.0 91.2
LLM prompting
Gemma v1.1 (7B) 69.7 42.0 71.7 70.9
Gemma v2 (9B) 83.0 39.7 88.5 67.5
Mistral v0.3 (7B) 74.2 31.2 75.9 70.9
Mixtral v0.1 (8x7B) 70.8 48.5 62.6 81.6
Llama v3.1 (8B) 81.8 33.6 84.5 63.3
Table 5: Precision and Recall scores for automatic evaluators on the QASemConsistency dataset. For most models, precision is substantially higher than recall.
Observing the F1 scores, the best F1 score on CLIFF is 62.2 and the best F1 score on FActScore is 76.6. In particular, as detailed in Table 5, precision tends to be significantly higher than recall, often with a substantial gap. For instance, Gemma 2 (9B) achieves a precision of 83.0 on CLIFF but a recall of only 39.7. This discrepancy suggests that LLMs may be over confident when generating a “Yes” response to indicate that a QA is supported. Hence, instead of relying solely on the model’s binary predictions, we suggest that users adjust the decision threshold in order to optimize performance.
6 Related Work
Identifying factual inconsistencies in attributable text generation has become a prominent research area in recent years. Existing methodologies for this task can be categorized based on the granularity of the detection. Table 1 illustrates and compares the various decomposition methods, recently proposed in the literature.
Starting with coarse-grained evaluation, SummEval (Fabbri et al., 2021) asks annotators to assign a 1-5 likert score to the entire summary, while some other works aim to produce a single score/label for the entire output (Yin et al., 2021; Rashkin et al., 2021; Honovich et al., 2022; Tang et al., 2023; Liu et al., 2023b). Several works evaluate each generated sentence separately with some (simple or sophisticated) form of aggregation (Falke et al., 2019; Kryscinski et al., 2020; Pagnoni et al., 2021; Utama et al., 2022; Tang et al., 2022; Laban et al., 2022; Gekhman et al., 2023; Mishra et al., 2024; Subbiah et al., 2024).
To achieve sub-sentence evaluation, a few recent works decompose each sentence into “claims”, “facts” or “propositions” (Table 1 Claim-level), whose support by the reference text is then assessed independently (Min et al., 2023; Krishna et al., 2023; Chen et al., 2023; Kamoi et al., 2023a; Wanner et al., 2024). Such decomposition has been also proposed for evaluating the summary content relevance, following the Pyramid paradigm (Nenkova and Passonneau, 2004; Liu et al., 2023c; Wan et al., 2024). However, these free text claims typically lack structure and systematicity. As a result, identifying unsupported claims does not effectively localize factual mistakes in the generated text. In fact, these claims are not atomic and encompass multiple semantic relations. For instance, the claim “The man died of measles” in Table 1 is not supported by the article, but it remains unclear whether the issue is that the man died for a reason other than measles, or that there is no man who died.
Another line of automatic evaluation methodology, typically referred to as question-generation and question-answering (QGQA), consists of generating questions and answers based on the model output and then comparing the answers to those obtained from the reference text by a QA model (Wang et al., 2020; Durmus et al., 2020; Nan et al., 2021; Scialom et al., 2021; Honovich et al., 2021; Fabbri et al., 2022). However, Kamoi et al. (2023b) demonstrate that this paradigm falls short in providing effective localization of factual inconsistency. This is primarily because the generated questions often contain factual inconsistencies from the summary itself. Indeed, as shown in Table 1 (QGQA), the generated question for the answer phrase “an inquest” assumes that the man died of measles and something was already opened and adjourned. In contrast, our method is based on predicate-argument level QAs, where each QA represents a single semantic relation.
Some other works ask human annotators to highlight inconsistent tokens or spans in the generated text (Maynez et al., 2020; Cao and Wang, 2021). This method often results in relatively poor inter-annotator agreement (e.g., 0.35 Fleiss Kappa in CLIFF), because span annotation is subjective (Mishra et al., 2024) and individual spans might serve multiple roles in the sentence, where some roles are in correct assertions and some are not. Indeed, the token-level evaluation in Table 1 includes the tokens “how he got the illness”, although this phrase also implies that the man got the illness, which is supported by the reference text. In that summary, the mistake is that the goal of the examination should be to determine the cause of death, rather than how the man contracted the illness. This unsupported fact cannot be captured with span highlighting while it can effectively identified using our QA “What didn’t something establish? how he got the illness” (Table 1). In the same line of work,  Laban et al. (2023) create SummEdits, a challenging benchmark with localized factual inconsistencies. However, unlike QASemConsistency, the mistakes are limited to a single token and are not naturally occurring.
Finally, Goyal and Durrett (2020) propose DAE, an automatic evaluation metric that operates at the level of semantic dependency arcs in a structured semantic dependency representation (Oepen et al., 2014) to localize inconsistent semantic relations in the generated text. Similarly, FactGraph (Ribeiro et al., 2022) represents both the source article and the summary with AMR (Banarescu et al., 2013) then model the factual inconsistencies at the edge level. However, these methods have only been applied automatically because obtaining human judgments at this level of granularity is challenging. Indeed, annotators would need to understand dependency or AMR labels and isolate the semantics of individual arcs within sentences, making manual evaluation difficult. Similarly to DAE and FactGraph, our QASemConsistency also operates at the level of individual semantic relations, while representing them with simple natural language question-answer pairs, enabling human evaluation as well and easing LLM processing. Finally, Cho et al. (2024) shown success at applying Neo-Davidsonian formal semantics to automatic text-to-image evaluation.
7 Limitations
Our work has several limitations. First, as we consider only verbal and nominal predicates, factual inconsistencies that stem from other predication types, e.g. adjectives, will not be localized with the finest granularity. For example, consider the reference text “John needs to repair his new red car, after the accident”, and the generated text “John’s blue car was damaged after he made an accident”. QASemConsistency would identify the inconsistency with the QA “What was damaged? John’s blue car”. However, this QA could be further divided into two smaller QAs “What was damaged? John’s car” and “What is blue? John’s car”, while only the latter is not supported. In addition, similarly to previous decomposition approaches (Min et al., 2023; Krishna et al., 2023), QASemConsistency does not capture factual inconsistencies that are due to implicit or inter-sentence discourse relations. For example, the sentences “John missed the train and arrived late” and “John arrived late and missed the train” would be treated as equivalent, even though they describe opposite causes and consequences. These limitations could be potentially addressed in future work by enriching QASemConsistency with additional semantic decompositions, such as QADiscourse (Pyatkin et al., 2020) or QA-Adj (Pesahov et al., 2023), when improved parsers will be made available.
8 Conclusion
We introduced QASemConsistency, a methodology that detects and localizes factual consistency errors to specific predicate-argument relations. Our error localization method is robust, as shown by our high human agreement. Contrast with other methods that either created complex, non-granular claims in natural language, or relied on linguistic formalisms that excluded non-expert annotators, our method can be applied with ease by non-expert users and models alike, without sacrificing granularity. Moreover, it can help human consumers of generative models recognize which parts of a response should be double checked. We hope that future work will focus on extending our approach to detect inconsistencies in the wider discourse, and help generative models to correct their responses.
Acknowledgments
We thank all the Mturk workers that participate in our annotation project. We thank Eran Hirsch, Aviv Slobodkin, Ayal Klein, Shuyang Cao, Vidhisha Balachandran, Tanya Goyal, Greg Durrett, Kyle Lo, Philippe Laban for their valuable discussions at various stages of the project. This research has been supported by a Google gift grant and the Israel Science Foundation (grant no. 2827/21). We also thank Google Cloud for providing us with credits for running experiments on the Google Cloud Platform. Arie Cattan is partially supported by the PBC fellowship for outstanding PhD candidates in data science.
Appendix A QASem parser
To automatically generate the QASem QAs for each generated text, we re-train the joint QASem parser Klein et al. (2022) and replace the original T5-small model with T5-XL (3B). This parser takes as input a sentence and a predicate, whether verbal or nominal, and generates a list of atomic QAs. We train our parser for 5 epochs until convergence with Adam optimizer and a learning rate of $055e-$ 05. We evaluate our parser on the QASRL gold data (Roit et al., 2020) and QANom (Klein et al., 2020), achieving 75.9 F1 (+7.3) on QASRL and 72.4 F1 (+13.2) on QANom. Similarly to (Klein et al., 2022), we use a span match threshold of IOU >= 0.3 to match between predicted and gold arguments.
Appendix B Human Evaluation
Filtering nonsensical QAs
Although our QASem parser significantly outperforms the current SOTA model (Klein et al., 2022) by a large margin (see §A), it is still prone to errors that can introduce noise into the consistency evaluation. Specifically, for nominal predicates, the parser may generate QAs that are nonsensical or convey a different meaning from the intended predicate. For instance, in the sentence “Sweden’s former foreign minister Johan Gustafssoon, who was kidnapped by Islamist militants in Mali in 2013, has been released after more than two years in captivity” (from CLIFF) the parser incorrectly generates ‘Where did someone militantise?” for the predicate “militants” and ‘Who was captained?” for the predicate “captivity”. In some rare cases, the parser hallucinates predicate argument relation. For instance, in the sentence “A man has been arrested on suspicion of murder after a man was found dead at a house in Cambridgeshire”, the parser generates the question “Where was someone arrested? at a house in Cambridgeshire”, while this location refers to the predicate “found”. To mitigate this issue, we add a preliminary annotation step, in which an annotator was asked to verify whether the QAs are both (1) semantically interpretable and (2) correctly represent the semantic relations in the generated text. This preliminary step is efficient and required less than an hour to validate 500 QAs.
Annotation Interface
Figure 3 shows the interface of the first annotation step (entity evaluation) and Figure 2 shows the interface of the second annotation step (QA verification). Once annotators finish the first step, they can move the second step, but they can always come back and modify their annotation for the entity evaluation.
Our annotation interface was developed with Vue.js [^10] https://vuejs.org as a WebComponent, which can embed into any website and popular annotation platforms, such as Amazon Mechanical Turk.
Figure 3: An example of the first annotation step.
Appendix C Automatic Evaluation
The prompt used in our experiments is shown in Table 6.
Instructions: In this task, you’ll be given an ARTICLE and an extremely simple question-answer (QA), representing a predicate-argument relation. Please indicate whether the QA is supported by the ARTICLE or not, according to the following definition: A QA is supported by the ARTICLE if the meaning of the QA can be deduced from the ARTICLE.
For example, given the article ‘‘Real Madrid beats PSG in the UEFA final Champions League’’: The QA ‘‘Who won something? PSG’’ is not supported because PSG didn’t win anything. The QA ‘‘What did someone win? The UEFA Champions League’’ is supported because someone (here Real Madrid) did win the UEFA Champions League.
ARTICLE:
QA:
Is the QA supported by the ARTICLE? Please answer only ‘‘Yes’’ or ‘‘No’’.
Table 6: Our prompt for automatically verifying whether a QASem QA is supported by the reference text, brown indicates an example.